You are a bidder in an independent private auction, and you value the object at $2000. Each bidder assumes that the valuations are uniformly distributed between $1000 and $5000. Determine your optimal bidding strategy in a first-price sealed bid auction when the total number of bidders are: 2, 10, and 100.
Answer #1

Total value of object = $2000

Lowest possible bidding = $1000

Highest possible bidding = $5000

Optimum bidding strategy

Optimum seal bid = Lowest possible valuation would be = $1000

Maximum value we can give = $2000

Maximum value - [ ( Maximum value - Lowest bid price ) No of bidders ]

When there are 2 Bidders = $2000 - [( $2000 - $1000 ) / 2] = $1500

When there are 10 Bidders = $2000 - [ ( $2000 - $1000) / 10 ] = $1900

  When there are 100 Bidders = $2000 - [ ( $2000 - $1000 ) / 100 ] = $1990